Circular No. 8107

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V4743 SAGITTARII S. Starrfield, Arizona State University; J.-U. Ness, Hamburg; J. J.

Drake, Harvard-Smithsonian Center for Astrophysics; and M. Orio, University of Wisconsin, report, for a much larger collaboration, that they observed V4743 Sgr for 7 hr beginning at Mar. 19.396 UT with the Chandra X-ray satellite (+ Low Energy Transmission Grating Spectrometer + High Resolution Camera): “The spectrum is complex, dominated by continuum emission from 2.0 to 4.0 nm (peak around 2.8 nm) with obvious and deep absorption features. Assuming a blue shift of −2400 km/s, we tentatively identify C VI at 3.38 nm and O VII at 2.16 nm and are in the process of identifying other spectral features. Our spectrum (qualitatively, but not quantitatively) resembles the XMM-Newton spectrum of Cal 83 by Paerels et al. (2001, A.Ap. 365, L308). It appears that we are observing the hot atmosphere of the underlying white dwarf undergoing nuclear burning near the surface. This nova is bright at x-ray wavelengths, with a count rate of 40 counts/s plus large amplitude oscillations of 10 counts/s with a period of 1323 s (preliminary). There appear to be additional periods present in the data, suggesting that this might be a pulsation rather than rotation, and further analysis of the light curve is ongoing (see also Drake et al. 2003, Ap.J. 584, 448). Of further note in the light curve, after 3.6 hr a slow decay occurs; the counts drop to near zero after ∼ 1.5 hr and remain near zero for the last 1.5 hr of the observation. We have a 36500-s XMM-Newton observation beginning at Apr. 4.921 and urge contemporaneous ground-based spectroscopic and photometric observations of this exciting nova.”

OBJECTS NEAR JUPITER V (AMALTHEA) P. D. Fieseler and S. M. Ardalan, Jet Propulsion Laboratory, California

Institute of Technology, report the discovery of several small objects south of Jupiter V in data obtained with the Galileo spacecraft’s star scanner on 2002 Nov. 5.26 UT during its closest approach. Assuming that the bodies are < 5 km in diameter, they all were within ∼ 5000 km of the spacecraft; limited data prevented the determination of more precise positions.

GRB 030329 Further to IAUC 8102, J. E. McGaha reports additional unfiltered

CCD magnitudes, calibrated to the R bandpass: Apr. 1.175 UT, 16.7; 1.184, 16.8; 1.208, 17.0; 1.241, 16.9; 1.259, 16.8; 1.261, 17.0; 3.162, 17.9; 1.179, 17.7; 3.188, 17.9; 3.199, 18.0; 4.170, 17.7; 4.191, 17.8; 4.199, 17.6.

2003 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green